OpenStudy (anonymous):

Suppose that the functions f and g are defined for all real numbers x as follows. f(x) = x + 2 g(x) = 4x + 3 Write the expressions for (f+g)(x) and (f-g)(x) and evaluate (f*g)(1).

OpenStudy (jchick):

Any idea how to start?

OpenStudy (anonymous):

I am not really sure, I kind of just need someone to walk me through this problem. I am having a hard time with it

OpenStudy (jchick):

Ok so have you done any so far?

OpenStudy (anonymous):

yes

OpenStudy (jchick):

Ok please post what you have done.

OpenStudy (anonymous):

Oh you meant on this problem?

OpenStudy (jchick):

Yes

OpenStudy (anonymous):

I haven't done anything on this problem. Sorry for the confusion

OpenStudy (jchick):

Ok so here is what we are looking at.

OpenStudy (anonymous):

So, x+2 + 4x + 3

OpenStudy (jchick):

We need to combine like terms!

OpenStudy (anonymous):

4x^2 + 5 ?

OpenStudy (jchick):

So combine the x and the constants.

OpenStudy (jchick):

Not quite look at 4x^2 again you are going to add one.

OpenStudy (anonymous):

Is it 4x^3 or add one to four

OpenStudy (jchick):

Add one to the four

OpenStudy (jchick):

So it will become 5x.

OpenStudy (jchick):

Can we combine our constants?

OpenStudy (anonymous):

That would be 5 as well

OpenStudy (jchick):

Yes so our answer is 5x + 5 would you like to try the other one?

OpenStudy (anonymous):

And that's the expression right?

OpenStudy (jchick):

We still have another one to solve for.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So x + 2 = 4x + 3

OpenStudy (anonymous):

Do we have to change the addition signs or does it stay the same?

OpenStudy (jchick):

All of the ones in the second will be minus.

OpenStudy (anonymous):

x - 2 - 4x -3?

OpenStudy (jchick):

I think I might have said it a little confusingly.

OpenStudy (jchick):

So now we need to combine like terms.

OpenStudy (anonymous):

-3x?

OpenStudy (anonymous):

for the x terms?

OpenStudy (jchick):

Yes that is correct

OpenStudy (jchick):

Now for the constants?

OpenStudy (anonymous):

2-3 = -1

OpenStudy (anonymous):

so -3x -1

OpenStudy (jchick):

Correct!

OpenStudy (jchick):

So we actually have one more thing to solve for.

OpenStudy (jchick):

Sorry didn't notice until now.

OpenStudy (anonymous):

great! okay so the next one

OpenStudy (jchick):

Ok so

OpenStudy (jchick):

Here is what we are looking at what is the one where the x usually is?

OpenStudy (jchick):

So what will we do with the 1?

OpenStudy (anonymous):

we multiply?

OpenStudy (jchick):

No we will use that instead of the x in (f)x

OpenStudy (jchick):

Sorry f(x)

OpenStudy (anonymous):

so f(1) = (4(1) + 3) * g(1) = (1 +2)

OpenStudy (anonymous):

g(1) = 8?

OpenStudy (anonymous):

or is it 7?

OpenStudy (anonymous):

and then 7*3 = 21

OpenStudy (jchick):

With PEMDAS it would be what?

OpenStudy (jchick):

Now we need to substitute it into the equation.

OpenStudy (anonymous):

Well I 4(1) + 3, which is 4(1) = 4 and then 4 +3 is 7?

OpenStudy (jchick):

Correct!

OpenStudy (anonymous):

so the answer is 21 right?

OpenStudy (anonymous):

3*7 = 21

OpenStudy (anonymous):

Is that right?

OpenStudy (jchick):

Yes!

OpenStudy (anonymous):

Thanks! Do you think you might be able to help me with one more question?

OpenStudy (jchick):

Sure!

OpenStudy (jchick):

By the way great job!

OpenStudy (anonymous):

Thank you! hold on one sec, I have to draw it out

OpenStudy (anonymous):

can you open this?

OpenStudy (anonymous):

try this one, sorry http://imgur.com/uByaXDx

OpenStudy (anonymous):

@jchick

OpenStudy (jchick):

Sorry I didn't get a reply for a few min so I left the page I will look.

OpenStudy (anonymous):

Okay!

OpenStudy (jchick):

I see the graph but where is the question?

OpenStudy (anonymous):

I am going to type it now! Use the graph of the function f below to find the following. All values at which f has a local minimum All local minimum values of f If there is more than one answer, separate them with commas

OpenStudy (anonymous):

@jchick

OpenStudy (jchick):

Ok sorry my computer wasn't letting me type anything.

OpenStudy (jchick):

Ok so they are looking for x. They want you to find the f(x).

OpenStudy (anonymous):

SO how do I do that?

OpenStudy (jchick):

Hang on I am drawing a blank here.

OpenStudy (anonymous):

This is what the question is looking for, all values at which f has a local minimum. do you think you can help me solve that first?

OpenStudy (jchick):

I believe that your local minimum is -3

OpenStudy (jchick):

You can find the local minimum by looking to see where the second to lowest point on the graph is.

OpenStudy (anonymous):

Ok so your right it's -3?

OpenStudy (anonymous):

what about the second question? how would i find that?

OpenStudy (anonymous):

all local minimum values of f?

OpenStudy (jchick):

What second equation?

OpenStudy (jchick):

Ok

OpenStudy (jchick):

I do not see F mentioned anywhere in your graph.

OpenStudy (jchick):

Oh wait I see it is a graph of F

OpenStudy (jchick):

You have an absolute minimum but I didn't know you could have more than one local minimum.

OpenStudy (anonymous):

Yeah this question confuses me :/

OpenStudy (jchick):

Hang on let me look at another resource.

OpenStudy (anonymous):

Okay